Mac's wholesale buys from a distributor which allows discounts of 25%, 20%, 5%.
A. What single discount ( called the single equivalent discount) would be equivalent to the above discounts?
B. Compare the above discounts to discounts of 30%, 20%.

Answer by ikleyn(53763)

.

(a)  (1-0.25)*(1-0.20)*(1-0.05) = 0.75*0.7*0.95 = 0.49875 = 1 - 0.50125.

     An equivalent single discount is 50.125%.



(b)  Two consecutive discounts of 30% and 20%  make the equivalent discount 

         (1-0.3)*(1-0.2) = 0.7*0.8 = 0.56 = 1 - 0.44

      of 44%.


     So, three consecutive discounts of (a) are "better" than two consecutive discounts of (b).

Solved and explained.

--------------

To see other numerous similar solved problems,  look into the lesson
    - Percentage problems on chains of discounts
in this site and have fun  (!)